Find x
log( 1/2, x ) = -1/3
The answer is 2^1/3
I want to know the method please.
Thank you

Answer by stanbon(75887)

Find x
log( 1/2, x ) = -1/3
The answer is 2^1/3
I want to know the method please.
------
log1/2(x) = -1/3
---
base = 1/2
log = exponent = -1/3
result = x
-----
x = (1/2)^(-1/3)
---
The negative in the exponent means "invert":
x = 2^(1/3)
==============
Cheers,
stan H.